Becker Office Service purchased a new computer system on January 1, 2018, for $38,900. It is expected to have a five-year useful life and a $3,700 salvage value. Becker Office Service expects to use the computer system more extensively in the early years of its life.

Required

  1. Calculate the depreciation expense for each of the five years, assuming the use of straight-line depreciation.
Year Cost Salvage Value Useful Life Annual Depreciation
1
2
3
4
5

b. Calculate the depreciation expense for each of the five years, assuming the use of double-declining-balance depreciation. (Enter all amounts as positive

values. Do not round intermediate calculations. Round "SL rate" answers to 2 decimal places. Round your answers to the nearest dollar amount.)

Accumulated Depreciation
Year Cost Beginning of Period (2 × SL Rate) Annual Depreciation
1
2
3
4
5
  1. Assume that Becker Office Service sold the computer system at the end of the fourth year for $19,000. Compute the amount of gain or loss using each depreciation method. (Negative amounts should be indicated with a minus sign. Do not round intermediate calculations. Round the final answers to nearest dollar amount.)
Amount Effect
Straight-Line
Double-Declining-Balance
Answer #1

Solution:

a) Calculation of Depreciation using Straight line depreciation method

Given Cost of computer = $38900

Salvage Value = $3700

Useful Life is 5 years

Depreciation = (cost - salvage value) / useful life

= (38900-3700)/5

= 35200/5

= $7040

In straight Line Depreciation method, Depreciation Remains same in every year i.e up to the useful life of asset

therefore deprecation for computer under straight line depreciation method is $7040 per year for five years.

Year

Opening Value ($)

Annual Depreciation

Closing Value ($)

(Opening Value – Depreciation)

1

38900

7040

31860

2

31860

7040

24820

3

24820

7040

17780

4

17780

7040

10740

5

10740

7040

3700

b) Calculation of Depreciation using Double Declining balance depreciation method

Depreciation Rate = 200%(100%/5)

= 200% * 20%

= 200/100 * 20%

= 2 * 20%

= 40%

Year

Opening Value ($)

Depreciation Rate

Annual Depreciation value ($)

Closing Value ($)

(Opening Value – Depreciation)

1

38900

40%

15560

23340

2

23340

40%

9336

14004

3

14004

40%

5601.6

8402.4

4

8402.4

40%

3360.96

5041.44

5

5041.44

40%

2016.58

3024.86

d) Calculating the amount of gain or loss under straight line depreciation method

Given Computer was sold at the end of Fourth year at $19000

Closing Value of computer at the end of fourth year is $10740

Profit = Sale value - Closing value

= 19000 - 10740

= $8260

Calculating the amount of gain or loss under Double Declining balance depreciation method

Closing Value of computer at the end of fourth year is $ 5041.44

Profit = Sale value - Closing value

= 19000 - 5041.44

= $13958.56. or $13959 (rounded off)

Note: if you have any doubts please comment for more explanation. Thank you